When a chemist titrates a standard solution of 308.25 mL of nitric acid (HNO3) with 0.155 M potassium hydroxide (KOH) , she finds that it requires 246.6 mL of the base to reach the endpoint of the titration. What is the molarity of the acid solution ?
1
Expert's answer
2020-04-24T13:34:51-0400

Solution:

The balanced chemical equation for the reaction between HNO3 and KOH:

HNO3 + KOH = KNO3 + H2O

At the endpoint of the titration: n(HNO3) = n(KOH)

C(HNO3) × V(HNO3) = C(KOH) × V(KOH);

C(HNO3) × 308.25 mL = (0.155 M × 246.6 mL);

C(HNO3) = (0.155 M × 246.6 mL) / (308.25 mL) = 0.124 M

Molarity of HNO3 = C(HNO3) = 0.124 M.


Answer: 0.124 M is the molarity of the acid solution (HNO3).